OpenStudy (anonymous):

y=(1/x)^ln x implicit differentation :(

OpenStudy (michele_laino):

is your expression as below: \[\Large y = {\left( {\frac{1}{x}} \right)^{\ln x}}\]

OpenStudy (anonymous):

\[y=(\frac{ 1 }{ x })^{\ln x}\]

OpenStudy (anonymous):

yes :)

OpenStudy (michele_laino):

ok! Then we have to take the logarithm (in base e) of both sides, so we get: \[\Large \ln y = \frac{{\ln x}}{x}\]

OpenStudy (anonymous):

won't be a minute :)

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

sorry about that :)

OpenStudy (michele_laino):

no worries :)

OpenStudy (alekos):

I got \[-(\ln x)^{2}\]

OpenStudy (michele_laino):

I'm sorry, I don't think that it is the right answer @alekos

OpenStudy (michele_laino):

now we have to compute the first derivative of both sides of the last equation. In order to that the first derivative of the left side, is given applying the chain rule, so we get: \[\Large \frac{{y'}}{y}\]

OpenStudy (michele_laino):

whereas in order to compute the first derivative of the right side, I apply the quotient rule, so I can write: \[\Large \left( {\frac{{\ln x}}{x}} \right)' = \frac{{\frac{1}{x}x - \ln x}}{{{x^2}}}\]

OpenStudy (michele_laino):

is it ok?

OpenStudy (anonymous):

yes got that :)

OpenStudy (michele_laino):

ok! so after the differentiation, we can write: \[\Large \frac{{y'}}{y} = \frac{{\frac{1}{x}x - \ln x}}{{{x^2}}}\]

OpenStudy (michele_laino):

now, please note that we can simplify the right side as below: \[\Large \frac{{\frac{1}{x}x - \ln x}}{{{x^2}}} = \frac{{1 - \ln x}}{{{x^2}}}\]

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

so we can rewrite that expression above, as below: \[\Large \frac{{y'}}{y} = \frac{{1 - \ln x}}{{{x^2}}}\]

OpenStudy (michele_laino):

now, in order to find the expresion of y', we have to multiply both sides by y, so what do you get?

OpenStudy (michele_laino):

expression*

zepdrix (zepdrix):

Michele I'm a little confused how you got this :o\[\Large \ln y = \frac{{\ln x}}{x}\]Wouldn't log of each side give us this?\[\Large\rm \ln y=\ln x\cdot \ln\left(\frac{1}{x}\right)\]

OpenStudy (anonymous):

\[y'=\frac{ y-\ln(xy) }{ x ^{2}y }\]

OpenStudy (michele_laino):

oops.. sorry you are right! @zepdrix I have made a big error sorry @Kerriy

OpenStudy (anonymous):

that is ok :)

zepdrix (zepdrix):

Ahhh darn, I wish we had caught that earlier hehe :)

OpenStudy (michele_laino):

:) @zepdrix

OpenStudy (michele_laino):

we have to restart, so the right expression is: \[\ln y = \ln x\ln \left( {\frac{1}{x}} \right) = \ln x \times \left( { - \ln x} \right) = - {\left( {\ln x} \right)^2}\]

OpenStudy (michele_laino):

so the first derivative of the right side is: \[ \Large - 2\ln x\left( {\frac{1}{x}} \right)\]

OpenStudy (michele_laino):

then we can write: \[\Large \frac{{y'}}{y} = - 2\ln x\left( {\frac{1}{x}} \right)\]

OpenStudy (anonymous):

where does the -lnx come from ?

OpenStudy (michele_laino):

since, using the properties of logarithm we have: \[\ln \left( {\frac{1}{x}} \right) = \ln 1 - \ln x = 0 - \ln x = - \ln x\]

OpenStudy (anonymous):

ok thank you so much :)

OpenStudy (michele_laino):

then in order to find the value of y' we have to multiply bot sides by y, so we get: \[\Large y' = - 2y\left( {\frac{1}{x}} \right)\ln x\]

OpenStudy (anonymous):

You are a champion, I cannot begin to thank you enough :) Can you tell me where I could go to learn more about logs and trig......or is it simply just practise?

OpenStudy (michele_laino):

you have to learn the basic properties, and then add much practise

OpenStudy (anonymous):

Thank you again, I really appreciate your effort and time :)

OpenStudy (michele_laino):

thank you! :) please in the last equation you ahve to replace y with its definition, namely: \[\Large y = {\left( {\frac{1}{x}} \right)^{\ln x}}\]

OpenStudy (michele_laino):

so your answer is: \[\Large y' = - 2{\left( {\frac{1}{x}} \right)^{\left( {\ln x + 1} \right)}}\ln x\]

OpenStudy (michele_laino):

Sorry again for my big error of before! @Kerriy

OpenStudy (anonymous):

It is fine you have been wonderful :)

OpenStudy (michele_laino):

Thanks! :)

OpenStudy (anonymous):

Welcome :)